OpenStudy (anonymous):

Please help! Use your function to determine how much you will owe the bank in five years. This is assuming you are not paying down your loan, so do not get scared! Show all of your work. Function: A(t) = 13,965 (1 + 03/12)^12t

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@confluxepic

OpenStudy (anonymous):

And this is what I have researched so far:

ganeshie8 (ganeshie8):

Alright! there is a mistake in your A(t) function. It should be : \[\large A(t) = 13,965 \left(1 + \dfrac{0.03}{12}\right)^{12t}\]

ganeshie8 (ganeshie8):

for #6, simply plugin \(t=5\) in above function and use your calculator to get a numerical answer

OpenStudy (anonymous):

Okay let me calculate that

OpenStudy (anonymous):

I got 839994.75

OpenStudy (anonymous):

Is that correct?

ganeshie8 (ganeshie8):

doesnt look correct, try again

ganeshie8 (ganeshie8):

you should get 16221.97.. http://www.wolframalpha.com/input/?i=13965%281+%2B+0.03%2F12%29%5E%2812*5%29

OpenStudy (anonymous):

Oh wait, no I plugged it in wrong I got 16221.97 as well

OpenStudy (anonymous):

Here are the last two questions

OpenStudy (anonymous):

Being a smart financial planner, you want to figure out how many months it will be until you owe the bank an amount of $100,000,00. Solve for t and show all your work

ganeshie8 (ganeshie8):

good :) for #7 you need to solve \(t\) in below equation : \[100,000 = 13,965 \left(1 + \dfrac{0.03}{12}\right)^{12t}\]

ganeshie8 (ganeshie8):

start by dividing \(13,965\) both sides

OpenStudy (anonymous):

7.16 = (1 + 0.03/12) ^12t

OpenStudy (anonymous):

Okay, what next?

ganeshie8 (ganeshie8):

Excellent! which is same as \(\large 7.16 = (1.0025)^{12t}\) taking \(\ln\) both sides you get \(\large \ln (7.16) = \ln(1.0025)^{12t} \) \(\large \ln (7.16) = 12t\ln(1.0025) \) \(\large t = \dfrac{\ln (7.16) }{12\ln(1.0025)}\)

ganeshie8 (ganeshie8):

use your calculator to get a numerical answer ^

ganeshie8 (ganeshie8):

thats it! we're done with #7

OpenStudy (anonymous):

So the answer for number 7 is 7.16 = (1.0025)^12t?

OpenStudy (anonymous):

Or do I need to solve further?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

But question number 8 is:

ganeshie8 (ganeshie8):

you need to solve further and find the time

ganeshie8 (ganeshie8):

first finish off #7

OpenStudy (anonymous):

So would I divide ln (7.16) by 12ln (1.0025)

ganeshie8 (ganeshie8):

Exactly! that gives the time in years, use your calculator

OpenStudy (anonymous):

I got 7.14 Years, not sure if that is correct though

ganeshie8 (ganeshie8):

wolfram says t=65.699 years http://www.wolframalpha.com/input/?i=ln+%287.16%29%2F%2812ln+%281.0025%29%29

ganeshie8 (ganeshie8):

multiply that by 12 to get the number of months it takes

OpenStudy (anonymous):

Ah I see now, thank you! Okay, last question

OpenStudy (anonymous):

Lastly, you decide to keep track of your loan four times a month instead of monthly. Solve for the adjusted interest rate. Remember to use the formula A(t)=P{(1+r/n)^1/c} where c = 4. When solving for the adjusted interest rate, be sure to set it equal to 1 + r/n

ganeshie8 (ganeshie8):

\[\large A(t) = 13,965 \left[1 + \dfrac{0.03}{12}\right]^{12t}\] rewrite it as \[\large A(t) = 13,965 \left[\left(1 + \dfrac{0.03}{12}\right)^{\frac{1}{4}}\right]^{4\times 12t}\] \[\large A(t) = 13,965 \left[1.00062\right]^{48t}\]

ganeshie8 (ganeshie8):

which is same as \[\large A(t) = 13,965 \left[1+0.00062\right]^{48t}\] so \(\dfrac{r}{48} = 0.00062\) solve \(r\) for the adjusted interest rate

OpenStudy (anonymous):

I got r = 0.02976

OpenStudy (anonymous):

@ganeshie8 would that be correct?